# Evolving Hard Maximum Cut Instances for Quantum Approximate Optimization Algorithms

### v1

@@ Abstract.
conference: Genetic and Evolutionary Computation Conference (GECCO ’25); July 14–18, 2025; Malaga, Spainbooktitle: Genetic and Evolutionary Computation Conference (GECCO ’25), July 14–18, 2025, Malaga, Spainprice: 15.00isbn: 978-1-4503-XXXX-X/18/06copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
Research has shown that certain algorithms underperform in specific scenarios, often labeled as hard or worst-case instances (Ko et al., 1986). Yet, guidance on how to generate or identify these challenging instances of studies is typically lacking in the literature. For instance, the RQAOA paper highlights an infinite family of $d$-regular bipartite graphs where the QAOA lags behind the GW algorithm when the depth (number of layers) is $O(log\textit{n})$, where n is the number of nodes (Bravyi et al., 2020). However, the paper does not provide instructions for constructing these graphs. In our work, we seek to bridge this gap by generating instances that underscore the differences between quantum algorithms and classical optimization algorithms in tackling the maximum cut problem, with a particular focus on the RQAOA and GW algorithms. While various heuristics have been explored in (Dunning et al., 2018), we selected the GW algorithm due to its well-established approximation bounds and extensive documentation of both its strengths and limitations, making it a suitable benchmark for comparison. We employ an evolutionary method using the classical Covariance Matrix Adaptation Evolution Strategy (CMA-ES) (Hansen and Ostermeier, 2001; Hansen et al., 2003) paired with a fitness function specifically tailored for this purpose. This strategy generates a variety of instances in the latent space of a thoroughly trained Graph Autoencoder (GAE). These instances are designed to be easily solvable by one algorithm while posing substantial challenges to the other, thereby providing a clearer comparison of their relative strengths and weaknesses.
@@
  • We have developed a method that effectively pinpoints maximum cut instances that either pose significant challenges or are considerably more manageable for the RQAOA in comparison to the GW algorithm.
@@ 2. PRELIMINARIES
crossing the two sets is maximal. The maximum cut problem is APX-hard (Papadimitriou and Yannakakis, 1991) and can be approximated using the GW algorithm (Goemans and Williamson, 1995). Specifically, this approach involves solving a relaxed version of the original problem. We first tackle an optimization problem in the continuous domain $[-1,1]^{N}$ using a semidefinite program, which provides us with a real-valued solution associated with a cost denoted as $C_{\text{relaxed}}$. Then, a sampling process, guided by probabilities derived from the entries of the real-valued solution (via a random projection method), is used to produce a feasible solution in the form of a binary string. It is important to note that the cost of the generated binary string is guaranteed to be at most equal to $C_{\text{relaxed}}$, and it is possible that $C_{\text{max}}<C_{\text{relaxed}}$. However, the expected value obtained by the cost function using this method is proven to be less than $\alpha C_{\text{max}}$, where $\alpha\approx 0.878$, for any input graph. In other words, this is an $\alpha$-approximation algorithm.
The maximum cut problem has received significant attention in the area of quantum optimization and different quantum algorithms have been designed. As discussed earlier, QAOA faces significant limitations due to its intrinsic locality and symmetry constraints, with substantial theoretical research having established no-go theorems in this area. To address these challenges, RQAOA (Bravyi et al., 2020) was introduced as an alternative quantum optimization algorithm specifically for combinatorial optimization problems. During each iteration of RQAOA, the problem size—represented as a graph or hypergraph—is reduced by one or more elements, based on correlations between variables derived from QAOA. This reduction process introduces non-local effects by creating new connections between nodes that were not previously linked, thus circumventing the locality constraints of QAOA. Finally, the iterative process stops once the number of variables is below a certain predefined threshold $n_{c}$ ($n_{c}\approx O(1)<<|V|$) such that a classical solver can be used to solve the remaining graph. In this study, we focus on depth-1 RQAOA, which can be classically simulated with a runtime of $O(n^{4})$ (Patel et al., 2024). Understanding the performance of these algorithms is challenging and pointing out instances where quantum approaches perform significantly better or worse than the classical GW algorithm provides an important understanding of their performance in comparison to state-of-the-art classical approximation algorithms.
Figure 1 presents the early performance of the RQAOA ($n_{c}$ is set to 10) over the first four iterations on a specifically generated 20-node maximum cut instance where it outperforms the GW algorithm. The diagram details the expectation values for each edge during the RQAOA process. In every iteration, the algorithm selects the edge that exhibits the highest absolute expectation value for operation. In cases where several edges possess identical maximum values, one edge is chosen at random. This sequential representation clearly illustrates the strategic process RQAOA employs to make decisions at each step.
Figure 1. Visualization of RQAOA’s performance ($n_{c}=10$) over the first four iterations on a 20-node maximum cut instance, where it outperforms the GW algorithm. Blue dots represent nodes, grey lines indicate edges, and the orange-highlighted node and edge denote the selected operation. The diagram tracks expectation values for each edge at every iteration, with RQAOA choosing the edge with the highest absolute expectation value, selecting randomly in case of a tie.
@@
GNNs have been proposed to tackle learning problems involving non-Euclidean data (Ju et al., 2024; Zhu et al., 2022), among which graph autoencoders (GAEs) are successfully applied to unsupervised tasks to learn graph topologies (Kipf and Welling, 2016). GAEs encode vertices or the entire graph into a latent space, where several different encoding/decoding methods are devised. GAEs have been used to extract network representations and learn the empirical distributions thereof (Wu et al., 2020). Due to the famous graph isomorphism issues, the generating/decoding component of GNNs is non-trivial: several works target this issue, including GraphRNN (You et al., 2018) employing a sequential method to construct nodes and edges by learning from a representative set of graphs; GraphVAE (Simonovsky and Komodakis, 2018) that uses variational autoencoders (Kingma and Welling, 2014) to generate graphs from continuous embeddings.
@@ 3. Related work
Evolutionary algorithms (EAs) have proven effective in generating diverse problem instances across various domains. Bossek et al. (2019) pioneered an EA-based approach for crafting the Travelling Salesperson Problem (TSP) instances using novel mutation operations. Neumann et al. (2019) developed a method for evolving diverse TSP instances and images using multi-objective indicators. Gao et al. (2021) introduced an EA-driven framework to generate problem instances classified as easy or hard based on distinct features. Addressing diversity in high-dimensional feature spaces, Marrero et al. (2023) proposed a Knapsack problem instance generator combining Principal Component Analysis with novelty search. More recently, research has explored evolving instances for the chance-constrained maximum coverage problem (Sadeghi Ahouei et al., 2024). This study departs from traditional approaches that rely on approximation ratios in the $(1+1)$ EA, instead incorporating performance ratio variances for improved evaluation.
@@ 4. Evolving Hard Maximum Cut Instances
GNNs have proven effective for generating graphs with diverse attributes. In our research, we use tools like the GAE to explore and identify specific graph structures. Our goal is to identify hard maximum cut instances. We begin by training a GAE with graph generators that are rooted in network science. After the GAE is adequately trained, we proceed to pinpoint particular graph instances within its complex latent space. The identification process is facilitated by the classical CMA-ES algorithm. This strategy employs a fitness function to assess scenarios where the RQAOA either surpasses or falls short compared to the GW algorithm.
@@
We choose to base our research on Permutation-Invariant Variational Autoencoder (PIGVAE) (Winter et al., 2021a) as it handles the graph isomorphism problem explicitly. The core principle of PIGVAE is that identical graphs should yield the same representation, regardless of the order of their nodes. It combines a variational autoencoder with a permuter model. This permuter model, trained in conjunction with the encoder and decoder, assigns a permutation matrix to each input graph. This matrix aligns the node order of the input graph with that of its reconstruction, ensuring consistent representation across different inputs.
@@
where $A$ is the baseline/reference algorithm and $B$ is the one of interest. In this paper, we try to answer two questions: is there a cluster/family of graphs (1) hard to solve by RQAOA w.r.t. the classical GW method; (2) vice versa? For those two questions, we maximize the objective functions, $\textit{p}_{\text{GW}}(\text{I})/\textit{p}_{\text{RQAOA}}$ and $\textit{p}_{\text{RQAOA}}(\text{I})/\textit{p}_{\text{GW}}$, respectively, in the latent space of the already-trained PIGVAE model. We utilize the widely applied CMA-ES for this maximization task. For each latent point sampled by the CMA-ES, we first decode the latent point to a graph instance $I$ and then execute RQAOA and GW algorithms on instance $I$ 100 times due to their stochasticity. We take the median value of each algorithm to compute the above ratio.
Importantly, a randomly sampled latent point may be decoded into a disconnected graph. We might initiate the CMA-ES with connected graphs. Still, as the search progresses, CMA-ES may sample disconnected graphs again. To resolve this issue, we propose the following strategy: for any latent point that corresponds to a disconnected graph, we solve the maximum cut problems supported on each connected component and then take the sum of the number of cuts from all components. For RQAOA, if the number of nodes of a component is smaller than a preset value $\textit{n}_{c}$, we solve the maximum cut problem on this component by a brute force method; Otherwise, we apply RQAOA on it. In the recursion step of RQAOA, disconnectedness can be created even if we start with a connected graph. In this case, we apply the above strategy to the connected components. In the experiments, we test our approach with two sizes: 20 nodes and 100 nodes, and we set $\textit{n}_{c}$ to 10 and 20, respectively.
@@ 5.1. Experimental setup
Figure 2. Second-largest eigenvalues of the normalized graph Laplacian for graph embeddings of 20-node (top) and 100-node (bottom) graphs. Nodes represent DBSCAN-identified clusters (Ester et al., 1996; Schubert et al., 2017), with edges indicating feature overlap or transitions between clusters. Node colors reflect the second-largest eigenvalue, offering insights into graph connectivity and structural complexity.
@@
We sampled the training graph instance u.a.r. from six random graph models: Erdős–Rényi (Erdos and Rényi, 1960), Watts-Strogatz small-world (Watts and Strogatz, 1998), Newman-Watts-Strogatz small-world (Newman and Watts, 1999), Random Regular (Steger and Wormald, 1999), Barabási–Albert (Barabási and Albert, 1999), and Dual-Barabási–Albert graphs (Moshiri, 2018). We also took the parameter setting of those graph models and hyperparameters of the PIGVAE model from (Winter et al., 2021b). In each training epoch, we generate 6 000 graph instances from the above random models. We train the PIGVAE model with maximally 3 000 epochs and a batch size of 128. An early stopping mechanism is implemented to terminate training if the validation loss shows no improvement over the past 60 epochs. To assess our approach on different graph sizes, we train two separate models: one for 20-node instances and the other for 100-node instances.
@@
The CMA-ES algorithm is configured with a population size of 64, leveraging parallel evaluations to enhance computational efficiency. To further optimize performance, a restart mechanism is implemented, allowing up to 10 restarts, with each limited to a maximum of 2,000 function evaluations. The algorithm is based on the standard implementation provided by the pycma library (Hansen et al., 2019).
@@
The full implementation is available in our Zenodo repository (Authors, s), along with supplementary materials for reproducibility and reference.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: Genetic and Evolutionary Computation Conference; July 14–18, 2025; Malaga, Spainbooktitle: Genetic and Evolutionary Computation Conference (GECCO ’25), July 14–18, 2025, Malaga, Spaindoi: 10.1145/3712256.3726415isbn: 979-8-4007-1465-8/2025/07ccs: Theory of computation Quantum computation theoryccs: Mathematics of computing Graph algorithmsccs: Computing methodologies Machine learningccs: Computing methodologies Evolutionary algorithmsccs: Theory of computation Approximation algorithms analysis
@@ 1. Introduction
Research has shown that certain algorithms underperform in specific scenarios, often labeled as hard or worst-case instances (Ko et al., 1986). Yet, guidance on generating or identifying such challenging instances remains limited in the literature. For instance, the RQAOA paper highlights an infinite family of $d$-regular bipartite graphs where the QAOA lags behind the GW algorithm when the depth (number of layers) is $O(log\textit{n})$, where n is the number of nodes (Bravyi et al., 2020). However, the paper does not provide concrete methods for constructing such graphs. In our work, we aim to bridge this gap by leveraging evolutionary computation method to generate instances that reveal the contrasting behaviors of quantum and classical optimization algorithms, specifically RQAOA and GW, when solving the maximum cut problem. While various heuristics have been explored in (Dunning et al., 2018), we select the GW algorithm due to its well-established approximation guarantees and extensive theoretical characterization, making it a suitable baseline for comparative analysis. From the evolutionary computation perspective, our contribution lies not in proposing a new variant of Covariance Matrix Adaptation Evolution Strategy (CMA-ES) (Hansen and Ostermeier, 2001; Hansen et al., 2003), but in repurposing it for a novel problem setting: rather than optimizing solutions to fixed problems, we use CMA-ES to evolve the problem instances themselves. Specifically, we operate within the latent space of a well-trained Graph Autoencoder (GAE), using CMA-ES to search for latent representations that yield graphs which are intentionally easy for one algorithm but hard for the other. The fitness function is carefully designed to reward maximum divergence in performance between RQAOA and GW.
@@
  • We introduce a method that effectively pinpoints maximum cut instances that either pose significant challenges or are considerably more manageable for the RQAOA in comparison to the GW algorithm.
@@ 2. PRELIMINARIES
crossing the two sets is maximal.
Goemans and Williamson algorithm
The maximum cut problem is APX-hard (Papadimitriou and Yannakakis, 1991) and can be approximated using the GW algorithm (Goemans and Williamson, 1995). Specifically, this approach involves solving a relaxed version of the original problem. We first tackle an optimization problem in the continuous domain $[-1,1]^{N}$ using a semidefinite program, which provides us with a real-valued solution associated with a cost denoted as $C_{\text{relaxed}}$. Then, a sampling process, guided by probabilities derived from the entries of the real-valued solution (via a random projection method), is used to produce a feasible solution in the form of a binary string. It is important to note that the cost of the generated binary string is guaranteed to be at most equal to $C_{\text{relaxed}}$, and it is possible that $C_{\text{max}}<C_{\text{relaxed}}$. However, the expected value obtained by the cost function using this method is proven to be less than $\alpha C_{\text{max}}$, where $\alpha\approx 0.878$, for any input graph. In other words, this is an $\alpha$-approximation algorithm.
Figure 1. RQAOA decision process ($n_{c}=10$) across the first four iterations on a 20-node maximum cut instance, where it surpasses the GW algorithm. Each subpanel represents one algorithm iteration, with nodes as blue dots and edges as grey lines. Edge labels show the expectation values derived from the quantum Hamiltonian (the quantum analog of the maximum cut cost function). At each step, the edge with the largest absolute expectation value is selected and contracted, highlighted in orange along with its endpoints, thus reducing the graph size by at least one node.
@@
Recursive Quantum Approximate Optimization Algorithm
The maximum cut problem has received significant attention in the area of quantum optimization and different quantum algorithms have been designed. As discussed earlier, QAOA faces significant limitations due to its intrinsic locality and symmetry constraints, with substantial theoretical research having established no-go theorems in this area. To address these challenges, RQAOA (Bravyi et al., 2020) was introduced as an alternative quantum optimization algorithm specifically for combinatorial optimization problems. During each iteration of RQAOA, the problem size—represented as a graph or hypergraph—is reduced by one or more elements, based on correlations between variables derived from QAOA. This reduction process introduces non-local effects by creating new connections between nodes that were not previously linked, thus circumventing the locality constraints of QAOA. Finally, the iterative process stops once the number of variables is below a certain predefined threshold $n_{c}$ ($n_{c}\approx O(1)\ll|V|$) such that a classical solver can be used to solve the remaining graph. In this study, we focus on depth-1 RQAOA, which can be classically simulated with a runtime of $O(n^{4})$ (Patel et al., 2024). Understanding the performance of these algorithms is challenging and pointing out instances where quantum approaches perform significantly better or worse than the classical GW algorithm provides an important understanding of their performance in comparison to state-of-the-art classical approximation algorithms.
Fig. 1 illustrates the initial decision-making process that RQAOA employs over four iterations, with $n_{c}=10$, on a generated 20-node maximum cut instance where RQAOA outperforms the GW algorithm. Each subfigure corresponds to one iteration, highlighting the single edge chosen for contraction. The selection is based on the largest absolute expectation value under the quantum state derived from QAOA: a positive sign indicates that the edge should remain uncut, while a negative sign indicates that it should be in the cut. In the event of a tie among these largest expectation values, any of the tied edges may be chosen arbitrarily. The selected edge and its corresponding nodes are highlighted in orange. The visualization offers an intuitive look at the way RQAOA navigates the solution space using quantum-informed decisions, distinguishing its behavior from classical algorithms. However, note that in this work we consider RQAOA at depth-1, which can be simulated classically, making it less clear whether its decisions might be parallel to those of some classical heuristic. However, at higher depths beyond classical simulability, RQAOA may have access to information not available to classical algorithms, offering a potential route to quantum advantage.
Features of Graphs
To effectively analyze the generated maximum cut instances, we compute a comprehensive suite of graph-level features that capture both intrinsic structural characteristics and algorithm-dependent behaviors. The structural features include classical topological indicators such as graph density, girth, chromatic number, assortativity, and transitivity, alongside spectral properties derived from the normalized Laplacian. Specifically, the spectral gap measures the difference between the top two eigenvalues and reflects how well-connected the graph is overall. The log-transformed largest and second-largest eigenvalues help identify structures like clusters or bottlenecks. Of particular interest is the log of the smallest non-zero eigenvalue, which indicates weak connectivity or the presence of nearly isolated nodes. We further include combinatorial descriptors like the normalized size of the maximum independent set to reflect sparsity and node connectivity. To evaluate how each instance interacts with the GW algorithm, we extract several features from its semidefinite programming (SDP) formulation, including the expected cut value, its variability across randomized rounding procedures, and statistics from the lower triangular portion of the SDP factor matrix. A complete list and detailed definitions of all features are presented in sections A and B (see the supplementary material (Pan et al., 2025)).
@@
GNNs have been proposed to tackle learning problems involving non-Euclidean data (Ju et al., 2024; Zhu et al., 2022), among which graph autoencoders (GAEs) are successfully applied to unsupervised tasks to learn graph topologies (Kipf and Welling, 2016). GAEs encode vertices or the entire graph into a latent space, where several different encoding/decoding methods are devised. GAEs have been used to extract network representations and learn the empirical distributions thereof (Wu et al., 2021). Due to the famous graph isomorphism issues, the generating/decoding component of GNNs is non-trivial: several works target this issue, including GraphRNN (You et al., 2018) employing a sequential method to construct nodes and edges by learning from a representative set of graphs; GraphVAE (Simonovsky and Komodakis, 2018) that uses variational autoencoders (Kingma and Welling, 2014) to generate graphs from continuous embeddings.
@@ 3. Related work
Evolutionary algorithms (EAs) have proven effective in generating diverse problem instances across various domains. Bossek et al. (2019) pioneered an EA-based approach for crafting the Travelling Salesperson Problem (TSP) instances using novel mutation operations. Neumann et al. (2019) developed a method for evolving diverse TSP instances and images using multi-objective indicators. Gao et al. (2021) introduced an EA-driven framework to generate problem instances classified as easy or hard based on distinct features. Addressing diversity in high-dimensional feature spaces, Marrero et al. (2023) proposed a Knapsack problem instance generator combining Principal Component Analysis with novelty search. More recently, research has explored evolving instances for the chance-constrained maximum coverage problem (Ahouei et al., 2024). This study departs from traditional approaches that rely on approximation ratios in the $(1+1)$ EA, instead incorporating performance ratio variances for improved evaluation.
[…]
Overview of the instance generation and evaluation pipeline.
Figure 2. Overview of the instance generation and evaluation pipeline.
@@ 4. Evolving Hard Maximum Cut Instances
GNNs have proven effective for generating graphs with diverse attributes. In our research, we use tools like the GAE to explore and identify specific graph structures. Our goal is to identify hard maximum cut instances. We begin by training a GAE with graph generators that are rooted in network science. After the GAE is adequately trained, we proceed to pinpoint particular graph instances within its complex latent space. The identification process is facilitated by the classical CMA-ES algorithm. This strategy employs a fitness function to assess scenarios where the RQAOA either surpasses or falls short compared to the GW algorithm. Furthermore, we perform a comprehensive analysis of the generated instances using a parallel coordinate plot and a dedicated machine learning pipeline. The overall method of the instance generation and evaluation pipeline is illustrated in the flowchart shown in Fig. 2.
@@
We choose to base our research on Permutation-Invariant Variational Autoencoder (PIGVAE) (Winter et al., 2021) as it handles the graph isomorphism problem explicitly. The core principle of PIGVAE is that identical graphs should yield the same representation, regardless of the order of their nodes. It combines a variational autoencoder with a permuter model. This permuter model, trained in conjunction with the encoder and decoder, assigns a permutation matrix to each input graph. This matrix aligns the node order of the input graph with that of its reconstruction, ensuring consistent representation across different inputs.
@@
where $A$ is the baseline/reference algorithm and $B$ is the one of interest. In this paper, we try to answer two questions: is there a cluster/family of graphs (1) hard to solve by RQAOA w.r.t. the classical GW method; (2) vice versa? For those two questions, we maximize the objective functions, $\textit{p}_{\text{GW}}(\text{I})/\textit{p}_{\text{RQAOA}}$ and $\textit{p}_{\text{RQAOA}}(\text{I})/\textit{p}_{\text{GW}}$, respectively, in the latent space of the already-trained PIGVAE model. We utilize the widely applied CMA-ES for this maximization task. For each latent point sampled by the CMA-ES, we first decode the latent point to a graph instance $I$ and then execute RQAOA and GW algorithms on instance $I$ 100 times due to their stochasticity. We take the median value of each algorithm to compute the above ratio. Importantly, a randomly sampled latent point may be decoded into a disconnected graph. We might initiate the CMA-ES with connected graphs. Still, as the search progresses, CMA-ES may sample disconnected graphs again. To resolve this issue, we propose the following strategy: for any latent point that corresponds to a disconnected graph, we solve the maximum cut problems supported on each connected component and then take the sum of the number of cuts from all components. For RQAOA, if the number of nodes of a component is smaller than a preset value $\textit{n}_{c}$, we solve the maximum cut problem on this component by a brute force method; Otherwise, we apply RQAOA on it. In the recursion step of RQAOA, disconnectedness can be created even if we start with a connected graph. In this case, we apply the above strategy to the connected components. In the experiments, we test our approach with two sizes: 20 nodes and 100 nodes, and we set $\textit{n}_{c}$ to 10 and 20, respectively.
Feature analysis
To better understand the structural factors that drive differences in algorithmic performance, we perform an in-depth feature analysis of the generated maximum cut instances. This analysis combines two complementary strategies: visualization through parallel coordinate plots and predictive modeling via a machine learning pipeline. We examine instances generated from multiple CMA-ES runs, focusing on both 20-node and 100-node graphs where RQAOA or the GW algorithm demonstrates superior performance. Each instance is described using 18 features, which include both general graph-theoretic properties and metrics derived from the GW algorithm’s semidefinite relaxation. The visualizations reveal characteristic trends in feature distributions tied to algorithmic success, while an AutoML-based classification framework, TPOT library (Olson et al., 2016), is used to train models capable of predicting the dominant algorithm for a given instance.
@@ 5.1. Experimental setup
Figure 3. Log-transformed second-largest eigenvalues of the normalized graph Laplacian for graph embeddings of 20-node (top) and 100-node (bottom) graphs. Nodes represent DBSCAN-identified clusters (Ester et al., 1996; Schubert et al., 2017), with edges indicating feature overlap or transitions between clusters. Node colors reflect the second-largest eigenvalue, offering insights into graph connectivity and structural complexity.
@@
We sample the training graph instance u.a.r. from six random graph models: Erdős–Rényi (Erdos and Rényi, 1960), Watts-Strogatz small-world (Watts and Strogatz, 1998), Newman-Watts-Strogatz small-world (Newman and Watts, 1999), Random Regular (Steger and Wormald, 1999), Barabási–Albert (Barabási and Albert, 1999), and Dual-Barabási–Albert graphs (Moshiri, 2018). We also take the parameter setting of those graph models and hyperparameters of the PIGVAE model from (Winter et al., 2021). In each training epoch, we generate 6 000 graph instances from the above random models. We train the PIGVAE model with maximally 3 000 epochs and a batch size of 128. An early stopping mechanism is implemented to terminate training if the validation loss shows no improvement over the past 60 epochs. To assess our approach on different graph sizes, we train two separate models: one for 20-node instances, trained using a single A100 GPU, and the other for 100-node instances, trained using four A100 GPUs.
@@
CMA-ES operates in the latent space of the GAE. The search space is bounded using per-dimension minima and maxima computed from the GAE’s training embeddings, ensuring the search remains within a region likely to yield valid graphs. The optimizer is configured with a population size of 64, leveraging parallel evaluations to enhance computational efficiency. To further optimize performance, a restart mechanism is implemented, allowing up to 10 restarts, with each limited to a maximum of 2 000 function evaluations. The algorithm is based on the standard implementation provided by the pycma library (Hansen et al., 2019).
Machine Learning classifier setup
To build classifiers, we label instances where RQAOA outperforms GW as 0, and those where GW outperforms RQAOA as 1. The dataset is split 70:30 into training and test sets. For the imbalanced 20-node dataset, we optimize for balanced accuracy (Brodersen et al., 2010), whereas standard accuracy (Pedregosa et al., 2011) is used for the balanced 100-node case. We also conduct 10-fold cross-validation over 100 generations, evaluating 60 candidate models per generation. To further investigate RQAOA’s reliability as a high-performance heuristic, we perform a secondary analysis using a modified labeling strategy. Here, instances are labeled as 1 if the ratio of GW to RQAOA performance is 0.96 or lower, indicating that RQAOA achieves at least a 4% improvement over GW. All other instances are labeled as 0.
@@
The full implementation is available in our Zenodo repository (Pan et al., 2025), along with supplementary materials for reproducibility and reference.